### v1

@@ 3.2 Two species dispersion relation
$\Lambda_{uv}(k)\equiv\text{cos}(\alpha_{uv})\,2\pi Up_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big{\rvert}_{V}\frac{\mu_{uv}}{\xi_{uv}^{2}}\,k\int_{0}^{\infty}\mathop{}\!\mathrm{d}s\,\Omega_{uv}\left(\frac{s}{\xi_{uv}}\right)s\,J_{1}(ks),$ (15)
@@
$\Lambda_{uv}(k)\equiv\text{cos}(\alpha_{uv})\,2\pi Up_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big{\rvert}_{V}\frac{\mu_{uv}}{\xi_{uv}}\,k^{2}\int_{0}^{\infty}\mathop{}\!\mathrm{d}s\,\Omega_{uv}\left(\frac{s}{\xi_{uv}}\right)s\,J_{0}(ks),$ (16)
by analogy with Equations (8) and (11), respectively. Both models share similar dispersion relations, with the only notable difference being the order of the Hankel transform of the interaction kernel, in their definitions of $\Lambda_{uv}$. The link between the two frameworks can be seen by redefining $\Omega_{uv}$ as the derivative of a different kernel $\omega_{uv}^{\prime}$ in Equation (15) and integrating by parts, which will yield Equation (16) but with the kernel $\omega_{uv}$. Crucially, in both equations, chirality only features through the cosine of the advection angle.
@@
$4\,\Lambda_{c\rho}\Lambda_{\rho c}+[(1-D)k^{2}+(\Lambda_{\rho\rho}-\Lambda_{cc})]^{2}\geq 0$ (18)
@@
3.2.1 Aside: assumptions about the kernel
To gain insight from these equations, we will assume that each $\Lambda_{uv}(k)$ shares the same sign as its corresponding $\mu_{uv}$ for values of $k$ that are ‘relevant for pattern formation’ – typically those corresponding to the largest positive values of $\lambda(k)$. From Equations (15) and (16), we see that this assumption requires that: (a) $p_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big{\rvert}_{V}>0$, which follows from our core model assumptions, and (b) the Hankel transforms of the interaction kernels are positive at these key values of $k$. We argue that this latter requirement is satisfied in many biologically relevant cases.
Condition (b) automatically holds for any kernel whose Hankel transform is always positive. Cho et al. [6] outline some sufficiency conditions for this. For both models, the primary condition is that $\sqrt{s}\Omega\left(\frac{s}{\xi}\right)$ is non-negative and decreasing. One way of satisfying this condition is by making a ‘splattergun’ assumption: as we increase distance, $s$, then in 2D space, interactions are spread out over a circle of increasingly large circumference $2\pi s$, and so we might expect the interaction kernel to decay at least as fast as $\frac{1}{s}$, analogously to how light intensity and some physical forces decay as $\frac{1}{s^{2}}$ in 3D space. For the gradient sensing model, Cho et al. [6] show that the $0^{\text{th}}$ order Hankel transform requires an additional condition for positivity, namely that $-\sqrt{s}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\Omega\left(\frac{s}{\xi}\right)$ is also non-negative and decreasing. Biologically speaking, all of the above conditions correspond to interactions that do not change direction with distance, and get sufficiently weaker as distance increases. Cho et al. [6] also stipulate that certain integrals of the kernel remain finite and that the kernel vanishes at infinity, both of which are naturally satisfied in biological contexts.
The above conditions are sufficient but not necessary, and many commonly studied kernels that do not satisfy them still have Hankel transforms that are always positive. In direct sensing models, for instance, both the exponential kernel, given by
$\Omega\left(\frac{s}{\xi}\right)=\frac{1}{2\pi}e^{-\frac{s}{\xi}},$ (19)
and the Gaussian-derivative kernel, $\Omega\left(\frac{s}{\xi}\right)\propto\frac{s}{\xi}e^{-0.5(\frac{s}{\xi})^{2}}$, used in Painter et al. [40], Jewell et al. [22], Salmaniw et al. [49], have positive $1^{\text{st}}$ order Hankel transforms. Similarly, for gradient sensing models, the exponential kernel and the Gaussian kernel, $\Omega\left(\frac{s}{\xi}\right)\propto e^{-0.5(\frac{s}{\xi})^{2}}$, used in Potts [46], Fagan et al. [11], Wang and Salmaniw [61], both have positive $0^{\text{th}}$ order Hankel transforms. Even in the case of the tophat kernel,
$\Omega\left(\frac{s}{\xi}\right)=\begin{cases}\frac{1}{\pi},\quad s\leq\xi\\ 0,\quad s>\xi\,,\end{cases}$ (20)
where the Hankel transform can be negative for some values of $k$, the values of $k$ for which this occurs often do not correspond to those with the highest values of $\lambda(k)$, as discussed in Jewell et al. [22]. Essentially, the negative values of the transform have smaller magnitude than the positive values, and also occur at higher values of $k$, where the diffusion term, $-k^{2}$, is larger, further reducing $\lambda$.
@@ 4.1 Direct sensing
Consider a single species distributed homogeneously in space with density $P$. Then, as shown in Fig. 2 a, consider a small perturbation in this density with amplitude $\delta$ that varies in space sinusoidally with wave-vector $\boldsymbol{k}$. If the perturbation grows, there is pattern formation. If it decays, the state returns to being spatially homogeneous. We can determine which outcome will occur by considering the advection induced at an arbitrary point, such as the red ‘$\times$’ in Fig. 2. Advection at ‘$\times$’ is induced by nonlocal interactions with agents at surrounding points. Consider points that are at some arbitrary distance, $s$, from ‘$\times$’. These are represented by the red circle in Fig. 2.
Now consider the dark yellow and blue points in Fig. 2, which are at a distance $s$ and are at an arbitrary angle $\theta$ and $-\theta$ relative to the direction of heterogeneity of the perturbation, $\boldsymbol{k}$. The advection induced at ‘$\times$’ by each point is shown as the arrow with corresponding colour in Fig. 2 b and c. The advection induced by both points will have the same magnitude, because there is an equal density of agents at both points. Thus, when advection is induced parallel to separation, we see that the overall net advection from both points is directed parallel to $\boldsymbol{k}$. In contrast, when advection is induced perpendicular to separation, the overall net advection from both points is directed perpendicular to $\boldsymbol{k}$. The angle $\theta$ is arbitrary, and so all points at distance $s$ (on the red circle) can be paired up in this way, such that the two points induce the same magnitude of advection and their net advection is parallel to $\boldsymbol{k}$ in the parallel case and perpendicular to $\boldsymbol{k}$ in the perpendicular case. The red ‘$\times$’ point and the distance ‘s’ are also arbitrary, so this argument applies to all points interacting at all distances.
@@
Figure 2: Schematic illustrating how the parallel component of a chiral interaction affects linear stability and pattern formation, whereas the perpendicular component does not. a A small sinusoidal perturbation of the density of a single-species with amplitude $\delta$ and wave-vector $\boldsymbol{k}$. We consider the advection at an arbitrary point (red ‘$\times$’) induced by interactions with points at distance $s$ (red circle). Points (dark yellow and blue) with opposite angles to the $\boldsymbol{k}$ direction are paired up, as they have the same density and so induce the same magnitude of advection. b Parallel interactions create a net advection parallel to the $\boldsymbol{k}$ direction, thereby changing the density distribution, affecting linear stability. c Perpendicular interactions create a net advection perpendicular to the $\boldsymbol{k}$ direction, and so only cause movement in directions that do not change the density distribution, and so do not affect linear stability.
@@ 4.2 Gradient sensing and assumptions
The argument presented above refers to the direct sensing model. Nevertheless, the conclusions also hold for the gradient sensing model. This can be shown directly, or by simply noting the relationship between the gradient and direct sensing models. For the former, consider the region within a circle around an arbitrary point, such as the red circle around the red ‘$\times$’ in Fig. 2. In the gradient sensing model, advection depends on the density gradient averaged across such a region. By symmetry, density only varies in the $\boldsymbol{k}$ direction, and so the gradient must point in this direction. Similarly to in the direct sensing model, advection that is induced parallel to this gradient is thus directed parallel to $\boldsymbol{k}$, affecting linear stability. Once again, advection induced perpendicular to the gradient is directed perpendicular to $\boldsymbol{k}$, and does not affect linear stability.
@@ 5 Insights on pattern formation
The linear stability analysis in Section 3 tells us that pattern formation occurs if $\text{Re}(\lambda_{+})>0$ for some $k>0$, where $\lambda_{+}(k)$ is specified by Equation (17). Additionally, we expect patterns to be stationary (rather than oscillating in time) if Condition (18) is satisfied, that is, if $\text{Im}(\lambda_{+}(k))=0$.
@@
Chirality presents a possible mechanism for suppressing these ‘detrimental’ effects of chase-and-run dynamics. If the runner or chaser species were to move more chirally, increasing the angle $\alpha_{\rho c}$ or $\alpha_{c\rho}$, then the magnitude of the $\Lambda_{c\rho}\Lambda_{\rho c}$ term would be reduced, which both promotes pattern formation and suppresses oscillations.
@@
Interestingly, a mechanism which acts in direct opposition to the effects of chase-and-run interactions is differential self-interaction. Differential self-interaction, where self-interactions between chasers differ in magnitude and/or direction to self-interactions between runners, is encoded in the term $[(1-D)k^{2}+(\Lambda_{\rho\rho}-\Lambda_{cc})]^{2}$ in Equations (17) and (18) (note that we include diffusion in this consideration of self-interaction). It is strictly positive, whereas chase-and-run terms, $\Lambda_{c\rho}\Lambda_{\rho c}$, are strictly negative. Differential self-interaction thus promotes pattern formation where chase-and-run suppresses it. Furthermore, the balance of these two effects determines whether emergent patterns are stationary or oscillatory (in the linear regime), as given by Condition (18). This condition depends only on the magnitude of the differential self-interaction, and not on whether it is the chaser or the runner that has the stronger interaction.
@@
Figure 3: Examples in which patterns fail to form with parallel chase-and-run dynamics but emerge when chase-and-run is chiral ($\alpha_{c\rho}=-20^{\circ}$ and $\alpha_{\rho c}=70^{\circ}$). This behaviour occurs across various model sensing types, kernels, and parameters (see Key). Dispersion curves show that the homogeneous steady-state is linearly stable for parallel chase-and-run (a) but unstable for chiral chase-and-run (b). Accordingly, simulations of cases in a remained homogeneous (not shown), while those of b formed patterns. Heatmaps of these patterns at late times are shown in c-h, with simulations run to $t=500$. Parameter values are in Table 1 of the Appendix. Heatmaps use the 2D colourmap in i, where high chaser density is yellow, high runner density is blue, and both species overlapping is green.
Figure 4: Examples in which patterns oscillate in time with parallel chase-and-run dynamics but are stationary when chase-and-run is chiral ($\alpha_{c\rho}=-20^{\circ}$ and $\alpha_{\rho c}=70^{\circ}$). This behaviour occurs across various model sensing types, kernels, and parameters (see Key). Dispersion curves show that the homogeneous steady-state is linearly unstable in all cases (a, b), but the imaginary part of $\lambda$ is non-zero for parallel chase-and-run (c) and zero for chiral chase-and-run (d). An example simulation of parallel chase-and-run (e) exhibits persistent oscillations, whereas the same system with chirality (f) reaches a stationary state that is stable for the remainder of the simulation. e and f use the colourmap from Fig. 3i, but in e, colours are rescaled for visibility ($c_{\text{min}},\,\rho_{\text{min}}=0.05$ and $c_{\text{max}},\,\rho_{\text{max}}=0.54$). Parameter values for each case (i–vi) are in Table 1 of the Appendix. Videos of these simulations are available at Jewell [21].
Figure 5: Simulation of parallel chase-and-run dynamics, initially exhibiting temporal oscillations before transitioning into a spot pattern. The spots then merge into a single larger spot for each species, reaching a state that is stationary at late times. In this case, linear theory predicts a complex growth rate ($\text{Im}(\lambda)\neq 0$), explaining the initial oscillation but failing to capture the long time behaviour. During the oscillatory and transition stages ($t=0$ to $t=41$), pattern amplitudes are small, $\mathcal{O}(10^{-3})$ to $\mathcal{O}(10^{-2})$, so colours from the colourmap in Fig. 3i are rescaled for visibility. After the pattern collapses into spots ($t=46$), amplitudes become large, $\mathcal{O}(10^{-1})$, and the standard $[0,1]^{2}$ colourmap range is used. Parameters match those of Fig.4 case (i), and are listed in Table 1 of the Appendix. A Video of this simulation is available at Jewell [21].
@@ 6.1 Choice of nonlinear functions
To test the predictions from our linear analysis, we simulate the full nonlinear integro-PDEs, namely Equation (5) for the direct sensing model, and Equation (6) for the gradient sensing model. We specify the volume-filling functions as $p_{c}(c,\rho)=p_{\rho}(c,\rho)=1-c-\rho$, and the $g_{uv}$ functions as $g_{cc}(c)=g_{\rho c}(c)=c$,  $g_{c\rho}(\rho)=g_{\rho\rho}(\rho)=\rho$. For the interaction kernels, we consider two commonly studied functions: the exponential kernel and the tophat kernel, defined in Equations (19) and (20), respectively.

### v2

@@ 3.2 Two species dispersion relation
$\Lambda_{uv}(k)\equiv\text{cos}(\alpha_{uv})\,2\pi U\phi_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big\rvert_{V}\frac{\mu_{uv}}{\xi_{uv}^{2}}\,k\int_{0}^{\infty}\Omega_{uv}\left(\frac{s}{\xi_{uv}}\right)s\,J_{1}(ks)\,\mathop{}\!\mathrm{d}s,$ (15)
@@
$\Lambda_{uv}(k)\equiv\text{cos}(\alpha_{uv})\,2\pi U\phi_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big\rvert_{V}\frac{\mu_{uv}}{\xi_{uv}}\,k^{2}\int_{0}^{\infty}\tilde{\Omega}_{uv}\left(\frac{s}{\xi_{uv}}\right)s\,J_{0}(ks)\,\mathop{}\!\mathrm{d}s,$ (16)
by analogy with Equations (8) and (11), respectively. Both models share similar dispersion relations, with the only notable difference being the order of the Hankel transform of the interaction kernel, in their definitions of $\Lambda_{uv}$. The link between the two frameworks can be seen by choosing $\Omega_{uv}=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\tilde{\Omega}_{uv}$ and integrating Equation (15) by parts, which will yield Equation (16). Crucially, in both equations, chirality only features through the cosine of the advection angle.
@@
$4\,\Lambda_{c\rho}\Lambda_{\rho c}+[(1-D)k^{2}+(\Lambda_{\rho\rho}-\Lambda_{cc})]^{2}\geq 0.$ (18)
@@
To gain insight from these equations, we will assume that each $\Lambda_{uv}(k)$ shares the same sign as $\mu_{uv}$ for values of $k$ that are ‘relevant for pattern formation’ – typically those corresponding to the largest positive values of $\lambda(k)$. From Equations (15) and (16), we see that this assumption requires that: (a) $\phi_{u}(C,P)\frac{\partial g_{uv}}{\partial v}\Big\rvert_{V}>0$, which follows from our core model assumptions, and (b) the Hankel transforms of the interaction kernels are positive at these key values of $k$. We argue in Appendix C that this latter requirement is satisfied in many biologically relevant cases.
@@ 4.1 Direct sensing
Consider a single species distributed homogeneously in space with density $P$. Then, as shown in Fig. 2 a, consider a small perturbation in this density with amplitude $\delta$ that varies in space sinusoidally with wave-vector $\boldsymbol{k}$. If the perturbation grows, there is pattern formation. If it decays, the state returns to being spatially homogeneous. We can determine which outcome will occur by considering the advection induced at an arbitrary point, such as the red $\times$’ in Fig. 2. Advection at $\times$’ is induced by nonlocal interactions with agents at surrounding points. Consider points that are at some arbitrary distance, $s$, from $\times$’. These are represented by the red circle in Fig. 2.
Now consider the dark yellow and blue points in Fig. 2, which are at a distance $s$ and are at an arbitrary angle $\theta$ and $\theta$ relative to the direction of heterogeneity of the perturbation, $\boldsymbol{k}$. The advection induced at $\times$’ by each point is shown as the arrow with corresponding colour in Fig. 2 b and c. The advection induced by both points will have the same magnitude, because there is an equal density of agents at both points. Thus, when advection is induced parallel to separation, we see that the overall net advection from both points is directed parallel to $\boldsymbol{k}$. In contrast, when advection is induced perpendicular to separation, the overall net advection from both points is directed perpendicular to $\boldsymbol{k}$. The angle $\theta$ is arbitrary, and so all points at distance $s$ (on the red circle) can be paired up in this way, such that the two points induce the same magnitude of advection and their net advection is parallel to $\boldsymbol{k}$ in the parallel case and perpendicular to $\boldsymbol{k}$ in the perpendicular case. The red $\times$’ point and the distance ‘s’ are also arbitrary, so this argument applies to all points interacting at all distances.
@@
Figure 2: Schematic illustrating how the parallel component of a chiral interaction affects linear stability and pattern formation, whereas the perpendicular component does not. a A small sinusoidal perturbation of the density of a single-species with amplitude $\delta$ and wave-vector $\boldsymbol{k}$. We consider the advection at an arbitrary point (red $\times$’) induced by interactions with points at distance $s$ (red circle). Points (dark yellow and blue) with opposite angles to the $\boldsymbol{k}$ direction are paired up, as they have the same density and so induce the same magnitude of advection. b Parallel interactions create a net advection parallel to the $\boldsymbol{k}$ direction, thereby changing the density distribution, affecting linear stability. c Perpendicular interactions create a net advection perpendicular to the $\boldsymbol{k}$ direction, and so only cause movement in directions that do not change the density distribution, and so do not affect linear stability.
@@ 4.2 Gradient sensing and assumptions
The argument presented above refers to the direct sensing model. Nevertheless, the conclusions also hold for the gradient sensing model. This can be shown directly, or by simply noting the relationship between the gradient and direct sensing models. For the former, consider the region within a circle around an arbitrary point, such as the red circle around the red $\times$’ in Fig. 2. In the gradient sensing model, advection depends on the density gradient averaged across such a region. By symmetry, density only varies in the $\boldsymbol{k}$ direction, and so the gradient must point in this direction. Similarly to in the direct sensing model, advection that is induced parallel to this gradient is thus directed parallel to $\boldsymbol{k}$, affecting linear stability. Once again, advection induced perpendicular to the gradient is directed perpendicular to $\boldsymbol{k}$, and does not affect linear stability.
@@ 5 Insights on pattern formation
Figure 3: Classification of the ($D$, $\alpha_{\rho c}$) parameter space in polar coordinates, based on the predictions of Equation (17), on a domain of size $L=8$. Increasing the magnitude of the chiral running angle, $\alpha_{\rho c}$, can shift the system from a linearly stable homogeneous state ($\text{Re}[\lambda(k)]<0$, blue) or a spatio-temporal instability ($\text{Re}[\lambda(k)]>0$, $\text{Im}[\lambda(k)]\neq 0$, red) to a stationary instability ($\text{Re}[\lambda(k)]>0$, $\text{Im}[\lambda(k)]=0$, green). Black $\circ,\times,\blacktriangleright$ markers indicate parameter values tested in numerical simulations and their patterning outcome. For this figure, $\alpha_{\rho\rho}=\alpha_{cc}=\alpha_{c\rho}=0$, with all other parameters given in Table 1 of Appendix A.
Before presenting numerical simulations, it is useful to first highlight the key conceptual insights from the linear stability analysis. This provides a framework for understanding how chirality and chase-and-run dynamics influence pattern formation and sets the stage for interpreting the subsequent full nonlinear dynamics. In particular, the linear stability analysis in Section 3 tells us that pattern formation occurs if $\text{Re}(\lambda_{+})>0$ for some $k>0$, where $\lambda_{+}(k)$ is specified by Equation (17). Additionally, we expect patterns to be stationary (rather than oscillating in time) if Condition (18) is satisfied, that is, if $\text{Im}(\lambda_{+}(k))=0$.
@@
Chirality presents a possible mechanism for suppressing these ‘detrimental’ effects of chase-and-run dynamics. If the runner or chaser species were to move more chirally, increasing the angle $\alpha_{\rho c}$ or $\alpha_{c\rho}$, then the magnitude of the $\Lambda_{c\rho}\Lambda_{\rho c}$ term would be reduced, which both promotes pattern formation and suppresses oscillations. This effect is illustrated in the phase diagram in Fig. 3, where increasing $\alpha_{\rho c}$ shifts the system out of regimes with either a stable homogeneous state or a spatio-temporal instability and into a regime with a stationary (non-temporal) instability.
@@
Interestingly, a mechanism which acts in direct opposition to the effects of chase-and-run interactions is differential self-interaction. Differential self-interaction, where self-interactions between chasers differ in magnitude and/or direction to self-interactions between runners, is encoded in the term $(1-D)k^{2}+(\Lambda_{\rho\rho}-\Lambda_{cc})]^{2}$ in Equations (17) and (18) (note that we include diffusion in this consideration of self-interaction). It is strictly positive, whereas chase-and-run terms, $\Lambda_{c\rho}\Lambda_{\rho c}$, are strictly negative. Differential self-interaction thus promotes pattern formation where chase-and-run suppresses it. Furthermore, the balance of these two effects determines whether emergent patterns are stationary or oscillatory (in the linear regime), as given by Condition (18). This condition depends only on the magnitude of the differential self-interaction, and not on whether it is the chaser or the runner that has the stronger interaction.
@@
Figure 4: Examples in which patterns fail to form with parallel chase-and-run dynamics but emerge when chase-and-run is chiral ($\alpha_{c\rho}=-20^{\circ}$ and $\alpha_{\rho c}=70^{\circ}$). This behaviour occurs across various model sensing types, kernels, and parameters (see Key). Dispersion curves show that the homogeneous steady-state is linearly stable for parallel chase-and-run (a) but unstable for chiral chase-and-run (b). Accordingly, simulations of cases in a remained homogeneous (not shown), while those of b formed patterns. Heatmaps of these patterns at late times are shown in c-h, with simulations run to $t=500$. Parameter values are in Table 1 of Appendix A. Heatmaps use the 2D colourmap in i, where high chaser density is yellow, high runner density is blue, and both species overlapping is green.
Figure 5: Examples in which patterns oscillate in time with parallel chase-and-run dynamics but are stationary when chase-and-run is chiral ($\alpha_{c\rho}=-20^{\circ}$ and $\alpha_{\rho c}=70^{\circ}$). This behaviour occurs across various model sensing types, kernels, and parameters (see Key). Dispersion curves show that the homogeneous steady-state is linearly unstable in all cases (a, b), but the imaginary part of $\lambda$ is non-zero for parallel chase-and-run (c) and zero for chiral chase-and-run (d). An example simulation of parallel chase-and-run (e) exhibits persistent oscillations, whereas the same system with chirality (f) reaches a stationary state that is stable for the remainder of the simulation. e and f use the colourmap from Fig. 4i, but in e, colours are rescaled for visibility ($c_{\text{min}},\,\rho_{\text{min}}=0.05$ and $c_{\text{max}},\,\rho_{\text{max}}=0.54$). Parameter values for each case (i–vi) are in Table 1 of Appendix A. Videos of these simulations are available at Jewell [21].
Figure 6: Simulation of parallel chase-and-run dynamics, initially exhibiting temporal oscillations before transitioning into a spot pattern. The spots then merge into a single larger spot for each species, reaching a state that is stationary at late times. In this case, linear theory predicts a complex growth rate ($\text{Im}(\lambda)\neq 0$), explaining the initial oscillation but failing to capture the long time behaviour. During the oscillatory and transition stages ($t=0$ to $t=41$), pattern amplitudes are small, $\mathcal{O}(10^{-3})$ to $\mathcal{O}(10^{-2})$, so colours from the colourmap in Fig. 4i are rescaled for visibility. After the pattern collapses into spots ($t=46$), amplitudes become large, $\mathcal{O}(10^{-1})$, and the standard $0,1]^{2}$ colourmap range is used. Parameters match those of Fig.5 case (i), and are listed in Table 1 of Appendix A. A Video of this simulation is available at Jewell [21].
@@ 6 Simulations of pattern formation
To test the predictions from our linear analysis, we simulate the full nonlinear integro-PDEs, namely Equation (5) for the direct sensing model, and Equation (6) for the gradient sensing model. These simulations allow us to assess how well the conceptual insights from the linear analysis extend into the fully nonlinear regime and to observe phenomena that arise beyond the scope of linear theory.
@@ 6.1 Choice of nonlinear functions
For all simulations in this study, we specify the volume-filling functions as $\phi_{c}(c,\rho)=\phi_{\rho}(c,\rho)=1-c-\rho$, and the $g_{uv}$ functions as $g_{cc}(c)=g_{\rho c}(c)=c$,  $g_{c\rho}(\rho)=g_{\rho\rho}(\rho)=\rho$. For the interaction kernels, we consider two commonly studied functions: the exponential kernel,
$\Omega\left(\frac{s}{\xi}\right)=\frac{1}{2\pi}e^{-\frac{s}{\xi}},$ (19)
and the tophat kernel,
$\Omega\left(\frac{s}{\xi}\right)=\begin{cases}\frac{1}{\pi},\quad s\leq\xi\\ 0,\quad s>\xi\,.\end{cases}$ (20)
In different examples, the interaction kernel is taken as either tophat or exponential, for both the direct sensing model ($\Omega$) and the gradient sensing model ($\tilde{\Omega}$). For simulations using the exponential kernel, the kernel’s spatial extent is cut-off at half the domain length ($\frac{L}{2}$) in all directions. This ensures that any two points only interact with each other once, rather than wrapping round the domain to interact from multiple directions. We have confirmed that this choice has a negligible impact on the linear stability analysis for all the examples in this study, as the domain length $L$ is always suitably larger than the decay scale of the exponential $\xi$.